OpenStudy (anonymous):

the averages of 5 consecutive integers starting with m as the first integer is n. what is the average of following integers? m+1,m+3,m+5,m+7,m+9

OpenStudy (perl):

[(m+1)+ (m+3)+ (m+5) + (m+7) + (m+9) ] / 5

OpenStudy (anonymous):

next step?

OpenStudy (perl):

then collect like terms

OpenStudy (bibby):

those are consecutive odds and evens. consecutive alone means 1,2,3,4,5.. (one right after the other) so here it'd be \(\large \frac{m+(m+1)+(m+2)+(m+3)+(m+4)}{5}=n\) they're asking for m+1,m+3,m+5,m+7,m+9 (which simplifies to 5m+25) our original fraction says m+m+1+m+2+m+3+m+4=5m+10 so we have the fact that \(\large \frac{5m+10}{5}=n\) and \(\large \frac{5m+25}{5}=n+\frac{15}{5}\)

OpenStudy (anonymous):

optiona are m+4 n+6 m+5

OpenStudy (anonymous):

@bibby

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

m+5? is ans i guess

OpenStudy (perl):

( m + 25 ) / 5 = m/5 + 5